# ELITR-Bench: A Meeting Assistant Benchmark for Long-Context Language Models

### v2

@@ 6 LLM-based evaluation assessment
In this section, we seek to verify the validity of the LLM-based (namely, GPT-4-based) evaluation methodology introduced in Section 4 and applied in Section 5. In Section 6.1, we define the LLM-based and human-based evaluators that we considered for comparison. The details of the crowdsourcing study we conducted to collect human score annotations are provided in Appendix C. Then, Section 6.2 presents our results and findings on the evaluator comparison.
@@ 6.1 Compared evaluators
  • Prometheus [21]: This fine-tuned model was originally proposed to provide an open-source alternative to using GPT-4 for score rubric-based evaluation. We used the Prometheus-13B-v1.0111111https://huggingface.co/kaist-ai/prometheus-13b-v1.0 model, with a prompt similar to the one adopted for GPT-4 – the only difference is that the score rubric is re-scaled to a 1-5 range to fit Prometheus’ expected format and multiplied by 2 in post-processing to be comparable to other scores. The prompt is available in Appendix E (Figs. 7 and 8).
  • Gold Human: This expert human annotation was done by one of the authors. The scores were assigned following the same 10-point score rubric as the one used for the GPT-4 evaluator (given in Appendix E, Fig. 6), to enforce consistency across questions.
  • Silver Human: This evaluator is based on a crowdsourcing study with the Prolific [^12] https://www.prolific.com/ platform where we averaged the scores assigned by 10 human annotators for each question. The annotators were provided with the same score rubric as for GPT-4 and Gold Human. We give more details on this evaluation in Appendix C.
Given the costly nature of human annotations, we performed our evaluation assessment on a small subset of the experiments described in Section 5.1. We specifically focused on the results of ELITR-Bench-QA’s test set in the single-turn mode. We looked at the results of 3 models that performed diversely in this setting: the best proprietary model (GPT-4), the best open-source model (Vicuna-13B-v1.5), and the worst open-source model (LongAlpaca-7B).
@@ 6.2 Evaluator comparison results
Table 4: Comparison of the scores obtained by different evaluators for the responses generated by GPT-4, Vicuna-13B-v1.5, or LongAlpaca-7B. The evaluation was performed on ELITR-Bench-QA’s test set in the single-turn mode, and for a single seeded run.
@@
To get a deeper understanding of how evaluators compare to one another, we calculated the Pearson correlation for every evaluator pair on the responses aggregated over the 8 meetings of the test set and generated by the three retained models. The results are displayed in Fig 2(b). GPT-4 shows a strong correlation with the two human-based evaluators (0.82 with Gold Human and 0.78 with Silver Human), which is in agreement with the findings from previous studies on GPT-4 judges [21, 4]. Prometheus, on the other hand, yielded a weak correlation (between 0.2 and 0.3) with all the other evaluators. We hypothesize that this could be due to a domain shift with respect to what Prometheus was fine-tuned on, caused by the nature of the meeting-related questions and the presence of anonymized entities (e.g., [PERSON3]). Turning to the two human-based evaluators, Gold Human and Silver Human obtained a very strong correlation of 0.89 which confirms the validity of the crowdsourcing study and the feasibility of the annotation task by non-expert judges.
@@
So far in this section, we have found that GPT-4 and human-based evaluators lead to scores that are highly correlated (see Figure 2(b)) but with slightly different score ranges (see Table 4). This led us to investigate how scores are distributed for different evaluators, and to study to what extent score levels match across evaluators. For that purpose, we considered the pool of (question, response, score) tuples obtained with the Gold Human evaluator on the responses from GPT-4, Vicuna-13B-v1.5, and LongAlpaca-7B for the 130 questions of the test set, i.e., 390 instances in total. We split these instances into 10 bins based on their score value from 1 to 10. Then, for all the instances in a bin, we check the distribution of the scores obtained by other evaluators on the bin’s (question, response) pairs. In practical terms, we seek to highlight through this procedure how Gold Human and alternative evaluators align at the grade level. The results are plotted in Fig. 2(a) where we describe the score distribution of the alternative evaluator through its means and 95% confidence intervals. Interestingly, we observe that the scores for the GPT-4 evaluator seem to fall into 3 distinct clusters, corresponding respectively to the intervals [1, 2], [3, 5] and [6, 10] in the Gold Human scores. This suggests that despite the use of a 10-point score rubric to align the GPT-4 evaluator’s scores with detailed desiderata, this evaluator is only able to distinguish between three levels of response quality. This finding then leads us to question the common practice of using LLM-based evaluator scores on a 5-point or 10-point scale. In contrast, the scores from Silver Human show a more linear relationship with the Gold Human scores, suggesting that implementing the 10-point score rubric in the crowdsourcing study aided in achieving a closer alignment between external human annotators and the evaluation criteria set by the organizers.
@@ 7 Conclusion
This paper introduced ELITR-Bench, a new benchmark for long-context LLMs focused on the meeting assistant task. We augmented the meeting transcripts from the existing ELITR corpus with 271 manually crafted questions and their respective ground-truth answers. Our experiments showcase the performance of recent long-context LLMs on ELITR-Bench, highlighting a gap between proprietary OpenAI models and LLaMA-2-based long-context models – in particular when dealing with multi-turn conversations. We validated our evaluation methodology based on a GPT-4 judge through its comparison against a Prometheus-based evaluator, as well as an expert human evaluator and a crowdsourcing-based evaluator. We demonstrate that the GPT-4 judge displays good correlation with human judgments, but a deeper investigation also reveals that it is unable to provide a very fine-grained evaluation on a 10-point scale, contrary to non-expert humans recruited on a crowdsourcing platform.
As future work, we are planning to extend ELITR-Bench for the evaluation of retrieval augmented generation (RAG) models. For instance, we could split each transcript into a set of short passages (containing a few utterances) and annotate the relevant passage(s) for each answer. Then, for each question, RAG models would need to first identify the relevant passage(s) and generate the response to the question based on those. Further studying the impact of de-identification (e.g., named entity anonymization) on QA performance is another interesting direction that could be investigated by reintroducing fake, randomly generated names into the meeting transcripts. In particular, we expect that this might have an impact on Who questions, which heavily depend on correctly generating anonymized entities that often have an important character overlap (e.g., [PERSON1] and [PERSON11]).
@@ Acknowledgments
The data collection and evaluation process rigorously adhered to the guidelines established by the UTTER EU project. In accordance with EU project policies, we regularly report to an ethics panel, with the most recent Ethical Review meeting held on November 2nd, 2023.
Notably, for the human evaluation of LLMs, we chose Prolific, an alternative crowdsourcing platform tailored for academic research. We meticulously followed Prolific’s guidelines for human experiments, deviating only in terms of compensation for human labelers. While Prolific sets a minimum compensation of \$6.50 per hour, we offered a significantly higher rate of £9 per hour (equivalent to \$11.5 per hour).
@@
The complete set of ELITR-Bench (question, ground-truth answer) pairs, along with metadata, is provided in JSON format at https://github.com/utter-project/ELITR-Bench. We also indicate for each question the responses generated by the different long-context LLMs considered in this paper, as well as the evaluation score attributed by the GPT-4 judge and other studied evaluators.
Additionally, the code for the response generation and for the evaluation will be released to enable reproducibility of the paper’s results and to foster future benchmarking efforts on ELITR-Bench.
Appendix A Experimental setup details
A.1 Compared models and hardware details
@@
We summarize the details of the different long-context LLMs compared in our experiments in Table 5. We provide for each model its context size limit in tokens, its backbone model (i.e., the pre-trained model used for the fine-tuning), and the link to the model checkpoint on Huggingface for open-source models or the link to the relevant OpenAI documentation for proprietary models.
The inference was done on a single A100 GPU with 80GB memory. In preliminary experiments, we also attempted to include the Mistral-7B-Instruct-v0.2131313https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 model in our study, as this model supports a context of up to 32K tokens. However, running this model on ELITR-Bench led to a GPU out-of-memory error on the A100, and thus we discarded it.
A.2 Configuration search on ELITR-Bench-QA’s dev set
In our pilot experiments, we noted that the open-source models retained for our study tended to be fairly impacted by the choice of the prompt and the inference hyperparameters. Therefore, we conducted a search on the inference configuration space to select appropriate hyperparameters for each open-source model. [^14] In comparison to open-source models, we found that GPT-4 and GPT-3.5 were more robust to differences in the inference configuration. Therefore, given the extensive cost of doing a large number of runs for commercial models, we did not conduct a configuration search on these. The configuration search was carried out in two steps on the dev set of ELITR-Bench-QA, in the single-turn mode. The evaluation was performed using GPT-4 as the evaluator, as described in the evaluation protocol in Section 4.
@@ A.2 Configuration search on ELITR-Bench-QA’s dev set
The specific chat template we adopted for each model is based on the one used during the model’s fine-tuning: the LLaMA2 template for LongAlpaca-7B and LongAlpaca-13B; the Vicuna template for LongChat-7B-v1.5, Vicuna-7B-v1.5 and Vicuna-13B-v1.5; and the LongAlign template for LongAlign-7B and LongAlign-13B.
@@
For the proprietary models, GPT-3.5 and GPT-4, we used nucleus sampling (temperature = 0.6 and top-p = 0.9) with the standard OpenAI chat template.
The cost of the two-step configuration search amounted to approximately \$150.161616We assessed the cost of performing the evaluation of a single model on the 141 dev set questions to \$3 approximately. As we evaluated 7 models on 6 configurations in the first step, and 7 models on 1 configuration in the second step, this yields \$147. To limit excessive expenses, we used the same model configuration for the different settings we experimented in (single-turn ELITR-Bench-QA, multi-turn ELITR-Bench-QA, and multi-turn ELITR-Bench-Conv).
@@
Appendix B Additional experimental results
B.1 Variance over seeded run results
@@
Note that the same seed is used both for the response generation part and the GPT-4-based evaluation part, as both can be sources of variance in the reported results. Based on our configuration search (see Appendix A.2), some of the response generation models were set to use greedy decoding: LongAlpaca-7B, LongAlpaca-13B, LongChat-7B-v1.5, LongAlign-7B. For such models, the response generation is deterministic and the only source of variance is that of the GPT-4 evaluator.
The results from Table 8 show that the variance across settings is fairly different. In the single-turn ELITR-Bench-QA setting, the standard deviation for all models remain relatively low, even for the models that use nucleus sampling (GPT-3.5, GPT-4, Vicuna-7B-v1.5, Vicuna-13B-v1.5, LongAlign-13B). However, in the multi-turn settings, we observe an increased standard deviation for those same models overall, in particular for Vicuna-7B-v1.5 and LongAlign-13B. We hypothesize that the sequence of questions asked in the same conversation in the multi-turn setting causes different seeded runs to cumulate errors and slightly diverge along the course of the conversation.
Table 8: Results for the seeded runs on the test set for different ELITR-Bench settings. The reported numbers correspond to the mean score ± sample standard deviation computed over 3 seeds. Boldface numbers correspond to the best performance among proprietary or open-source models. The results for GPT-3.5 are omitted in the multi-turn setting as the context length exceeded the 16K limit of this model.
@@
B.2 Additional results on question type and answer position
Table 9: Results by question type and answer position on the test set of ELITR-Bench-QA in single-turn mode. The number N below a subset indicates the corresponding subset size.
Model Question type Answer position
Who
(N=45)
What
(N=57)
When
(N=20)
How many
(N=8)
Begin
(N=43)
Middle
(N=34)
End
(N=22)
Several
(N=31)
GPT-3.5 7.91 6.94 7.68 7.79 7.33 7.45 7.76 7.37
GPT-4 8.56 8.29 8.28 8.29 8.36 8.29 8.32 8.57
LongAlpaca-7B 5.35 5.37 6.35 6.79 5.81 5.80 4.97 5.53
LongAlpaca-13B 7.19 5.47 6.47 6.00 5.93 5.95 6.85 6.59
LongChat-7B-v1.5 6.88 4.94 6.33 4.17 6.41 4.91 5.89 5.77
Vicuna-7B-v1.5 6.13 5.65 5.40 2.88 5.89 5.21 4.96 6.12
Vicuna-13B-v1.5 6.96 6.68 5.48 5.54 6.35 6.41 6.55 6.87
LongAlign-7B-64k 6.93 6.33 6.00 5.88 7.09 6.39 6.47 5.66
LongAlign-13B-64k 6.08 6.74 5.97 5.75 6.71 6.21 6.33 5.95

### v3

@@ 6 LLM-based evaluation assessment
In this section, we seek to verify the validity of the LLM-based (namely, GPT-4-based) evaluation methodology introduced in Section 4 and applied in Section 5. In Section 6.1, we define the LLM-based and human-based evaluators that we considered for comparison. The details of the crowdsourcing study we conducted to collect human score annotations are provided in Appendix D. Then, Section 6.2 presents our results and findings on the evaluator comparison.
@@ 6.1 Compared evaluators
  • Prometheus [23]: This fine-tuned model was originally proposed to provide an open-source alternative to using GPT-4 for score rubric-based evaluation. We used the Prometheus-13B-v1.0131313https://huggingface.co/kaist-ai/prometheus-13b-v1.0 model, with a prompt similar to the one adopted for GPT-4 – the only difference is that the score rubric is re-scaled to a 1-5 range to fit Prometheus’ expected format and multiplied by 2 in post-processing to be comparable to other scores. The prompt and the score rubric are detailed in Appendix F (Figs. 8 and 9, respectively).
  • Gold Human: This expert human annotation was done by one of the authors. The scores were assigned following the same 10-point score rubric as the one used for the GPT-4 evaluator (given in Appendix F, Fig. 7), to enforce consistency across questions.
  • Silver Human: This evaluator is based on a crowdsourcing study with the Prolific [^14] https://www.prolific.com/ platform where we averaged the scores assigned by 10 human annotators for each question. The annotators were provided with the same score rubric as for GPT-4 and Gold Human. We give more details on this evaluation in Appendix D.
Given the costly nature of human annotations, we performed our evaluation assessment on a small subset of the experiments described in Section 5.1. We specifically focused on the results of ELITR-Bench-QA’s test set in the single-turn mode. We looked at the results of 3 models that performed diversely in this setting: GPT-4, Vicuna-13B-v1.5, and LongAlpaca-7B. [^15] The crowdsourcing study was conducted in February 2024 – i.e., before the release of Phi-3 and LLaMA-3.1 – which is why these more recent models were not included in the evaluation assessment experiments.
@@ 6.2 Evaluator comparison results
Table 4: Comparison of the scores obtained by different evaluators for the responses generated by GPT-4, Vicuna-13B-v1.5, or LongAlpaca-7B. The reported numbers correspond to the average scores from 1 to 10 (higher is better) obtained on ELITR-Bench-QA’s test set in the single-turn mode, and for a single seeded run.
@@
To get a deeper understanding of how evaluators compare to one another, we calculated the Pearson correlation for every evaluator pair on the responses aggregated over the 8 meetings of the test set and generated by the three retained models. The results are displayed in Fig. 2(b). GPT-4 shows a strong correlation with the two human-based evaluators (0.82 with Gold Human and 0.78 with Silver Human), which is in agreement with the findings from previous studies on GPT-4 judges [23, 5]. Prometheus, on the other hand, yielded a weak correlation (between 0.2 and 0.3) with all the other evaluators. We hypothesize that this could be due to a domain shift with respect to what Prometheus was fine-tuned on, caused by the nature of the meeting-related questions and the presence of anonymized entities (e.g., [PERSON3]). Turning to the two human-based evaluators, Gold Human and Silver Human obtained a very strong correlation of 0.89 which confirms the validity of the crowdsourcing study and the feasibility of the annotation task by non-expert judges.
@@
So far in this section, we have found that GPT-4 and human-based evaluators lead to scores that are highly correlated (see Fig. 2(b)) but with slightly different score ranges (see Table 4). This led us to investigate how scores are distributed for different evaluators, and to study to what extent score levels match across evaluators. For that purpose, we considered the pool of (question, response, score) tuples obtained with the Gold Human evaluator on the responses from GPT-4, Vicuna-13B-v1.5, and LongAlpaca-7B for the 130 questions of the test set, i.e., 390 instances in total. We split these instances into 10 bins based on their score value from 1 to 10. Then, for all the instances in a bin, we check the distribution of the scores obtained by other evaluators on the bin’s (question, response) pairs. In practical terms, we seek to highlight through this procedure how Gold Human and alternative evaluators align at the grade level. The results are plotted in Fig. 2(a) where we describe the score distribution of the alternative evaluator through its means and 95% confidence intervals. Interestingly, we observe that the scores for the GPT-4 evaluator seem to fall into 3 distinct clusters, corresponding respectively to the intervals [1, 2], [3, 5] and [6, 10] in the Gold Human scores. This suggests that despite the use of a 10-point score rubric to align the GPT-4 evaluator’s scores with detailed desiderata, this evaluator is only able to distinguish between three levels of response quality. This finding then leads us to question the common practice of using LLM-based evaluator scores on a 5-point or 10-point scale. In contrast, the scores from Silver Human show a more linear relationship with the Gold Human scores, suggesting that implementing the 10-point score rubric in the crowdsourcing study aided in achieving a closer alignment between external human annotators and the evaluation criteria set by the organizers.
@@ 7 Conclusion
This paper introduced ELITR-Bench, a new benchmark for long-context LLMs focused on the meeting assistant task. We augmented the meeting transcripts from the existing ELITR corpus with 271 manually crafted questions and their respective ground-truth answers. We also produced noisy versions of the transcripts to study the impact of noise on models’ performance. Our experiments confirmed the improvements achieved through successive model iterations (e.g., GPT-3.5 vs GPT-4/4o, or LLaMA-2 vs LLaMA-3.1). We also found that while the best tested open models Phi-3 and LLaMA-3.1 approach the performance of OpenAI’s frontier model GPT-4o on clean transcripts, there remains a gap in noise robustness between these models. We validated our evaluation methodology based on a GPT-4 judge through its comparison against a Prometheus-based evaluator, as well as an expert human evaluator and a crowdsourcing-based evaluator. We demonstrate that the GPT-4 judge displays good correlation with human judgments, but a deeper investigation also reveals that it is unable to provide a very fine-grained evaluation on a 10-point scale, contrary to non-expert humans recruited on a crowdsourcing platform.
As future work, we are considering extending ELITR-Bench for the evaluation of retrieval augmented generation (RAG) models. For instance, we could split each transcript into a set of short passages (containing a few utterances) and annotate the relevant passage(s) for each answer. Then, RAG models would generate the response to a question using the retrieved passage(s). Further studying the impact of de-identification (e.g., named entity anonymization) on QA performance is another interesting direction that could be investigated by reintroducing fake, randomly generated names into the meeting transcripts. We expect that this might have an impact on Who questions, which heavily depend on correctly generating anonymized entities that often have an important character overlap (e.g., [PERSON1] and [PERSON11]).
@@ Acknowledgments
The data collection and evaluation process rigorously adhered to the guidelines established by the UTTER EU project. In accordance with EU project policies, we regularly report to an ethics panel, with the most recent Ethical Review meeting held on September 5th, 2024.
Notably, for the human evaluation of LLMs, we chose Prolific, a crowdsourcing platform tailored for academic research. We meticulously followed Prolific’s guidelines for human experiments, deviating only in terms of compensation for human labelers. While Prolific sets a minimum compensation of \$6.50 per hour, we offered a significantly higher rate of £9 per hour (equivalent to \$11.5 per hour).
@@
The complete set of ELITR-Bench (question, ground-truth answer) pairs, along with metadata, is publicly available at https://github.com/utter-project/ELITR-Bench. We also indicate for each question the responses generated by the different long-context LLMs considered in this paper, as well as the evaluation score attributed by the GPT-4 judge and other evaluators studied.
Additionally, the code for the response generation and for the evaluation was released alongside the data to enable reproducibility of the paper’s results and to foster future benchmarking efforts on ELITR-Bench.
Appendix A Noise injection in meeting transcripts
To evaluate the robustness of long-context models to noisy text, we generated multiple noisy versions of the ELITR meeting transcripts by simulating various levels of automatic speech recognition (ASR) noise. We utilized a large corpus of over 500,000 ASR transcripts aligned with reference texts, derived from the LibriSpeech corpus [38] and decoded using the Google Cloud Speech-to-Text API. The dataset includes annotated transcription errors and is available from the RED-ACE ASR Error Detection and Correction dataset. [^16] https://huggingface.co/datasets/google/red_ace_asr_error_detection_and_correction We combined the train, development, and test sets into a single JSON file with 525,308 entries, using it to generate 86,148 substitution rules by adapting code from an ASR error simulator. [^17] https://github.com/pe-trik/asr-errors-simulator Each rule specifies a token and a probability distribution over similar tokens (or an empty character) that can replace it. An example of a such a substitution rule is as follows:
"earlier": {
"early": 0.4651, "Elliot": 0.0233, "Early": 0.0233,
"Hurley": 0.0233, "earliest": 0.1860, "other": 0.0465,
"earrings": 0.0233, "early.": 0.0233, "area": 0.0233,
"earlier.": 0.0698, "Julia": 0.0233, "audio": 0.0233,
"Aaliyah": 0.0233, "Italia": 0.0233
These rules enable the simulation of noisy transcripts from existing ‘clean’ transcripts, targeting specific Word Error Rate (WER) levels of 20%, 40%, 60%, 80%, and 100%. However, the actual WER achieved after noise injection is generally lower than the intended target, meaning that even at a target WER of 100%, some correct answers can still be inferred. The actual WER values corresponding to different target levels are presented in Table 5.
Table 5: Comparison of target Word Error Rates with the effective Word Error Rates obtained by the noise injection procedure, averaged over all dev and test meeting transcripts.
Target WER (%) Effective WER (%)
20.00 20.99
40.00 37.41
60.00 50.07
80.00 61.86
100.00 71.13
Below is an example of GPT-4’s response to the same question when provided with transcripts from the same meeting but with varying levels of noise in the long-context window. The answers below correspond to the question “Who is going to register for the [PROJECT2] conference?”:
  • Ground truth: “[PERSON3]”
  • 20% WER: “(PERSON3) is going to register for the [PROJECT2] conference.”
  • 40% WER: “The transcript does not explicitly state who will register for the [PROJECT2] conference.”
  • 60% WER: “Based on the transcript, (PERSON3) indicated that they can register for the [PROJECT2] conference.”
  • 80% WER: “Based on the transcript, it is not entirely clear who specifically is going to register for the [PROJECT2] conference.”
We observe that GPT-4 remains robust to text noise across several noise levels for this query. It is important to note that a poor response at 40% WER and a better response at higher noise levels can occur because the target WERs (noise levels) are calculated for the entire meeting transcript. Various segments of the transcript, especially the relevant ones, may be impacted in diverse ways or remain unaffected at different noise levels.
Appendix B Experimental setup details
B.1 Compared models and hardware details
@@
LLaMA-3.1-8B 128k LLaMA-3.1-8B https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct
Phi-3-small 128k Phi-3-small https://huggingface.co/microsoft/Phi-3-small-128k-instruct
In our experiments on ELITR-Bench, we compared responses generated by 12 LLMs with long-context capabilities. We included both commercial models and open long-context models in our benchmarking:
  • GPT-3.5, GPT-4 [36] and GPT-4o [^18] https://openai.com/index/gpt-4o-system-card/ are powerful commercial LLMs from OpenAI that have obtained state-of-the-art performance on a wide range of LLM benchmarks. We used the gpt-3.5-turbo-16k-0613, gpt-4-1106-preview, and gpt-4o-2024-05-13 checkpoints, [^19] https://platform.openai.com/docs/models/ which enable a context length of 16k tokens for GPT-3.5 and 128k for GPT-4 and GPT-4o.
  • LLaMA-3.1-8B is the latest iteration (at the time of writing) of the LLaMA family of models from Meta AI [17]. This model enables a context limit of 128k due to its native long-context fine-tuning. We used the instruction-tuned version of the model in our experiments.
  • Phi-3-small is the 3rd model of the Phi family of LLMs from Microsoft [1]. We adopted the ‘small’ instruction-tuned version which has 7B parameters and accepts a context up to 128k tokens.
We summarize the details of the different long-context LLMs in Table 6. We provide for each model its context size limit in tokens, its backbone model (i.e., the pre-trained model used for the fine-tuning), and the link to the model checkpoint on Huggingface for open models or the link to the relevant OpenAI documentation for proprietary models.
The inference was done on a single A100 GPU with 80GB memory. In preliminary experiments, we also attempted to include the Mistral-7B-Instruct-v0.2202020https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 model in our study, as this model supports a context of up to 32k tokens. However, running this model on ELITR-Bench led to a GPU out-of-memory error on the A100, and thus we discarded it.
B.2 Configuration search and hyperparameter setting
In our pilot experiments, we noted that the LLaMA-2-based open models retained for our study tended to be fairly impacted by the choice of the prompt and the inference hyperparameters. Therefore, we conducted a search on the inference configuration space to select appropriate hyperparameters for each LLaMA-2-based model. [^21] In comparison to these, we found that GPT-3.5, GPT-4/4o, LLaMA-3.1-8B, and Phi-3-small were more robust to differences in the inference configuration. Therefore, we did not conduct a configuration search on these. The configuration search was carried out in two steps on the dev set of ELITR-Bench-QA, in the single-turn mode. The evaluation was performed using GPT-4 as the evaluator, as described in the evaluation protocol in Section 4.
@@ B.2 Configuration search and hyperparameter setting
The specific chat template we adopted for each model is based on the one used during the model’s fine-tuning: the LLaMA-2 template for LongAlpaca-7B and LongAlpaca-13B; the Vicuna template for LongChat-7B-v1.5, Vicuna-7B-v1.5 and Vicuna-13B-v1.5; and the LongAlign template for LongAlign-7B and LongAlign-13B.
@@
The cost of the two-step configuration search amounted to approximately \$150.232323We assessed the cost of performing the evaluation of a single model on the 141 dev set questions to \$3 approximately. As we evaluated 7 models on 6 configurations in the first step, and 7 models on 1 configuration in the second step, this yields \$147. To limit excessive expenses, we used the same model configuration for the different settings we experimented in (single-turn ELITR-Bench-QA, multi-turn ELITR-Bench-QA, and multi-turn ELITR-Bench-Conv).
For the proprietary models, GPT-3.5, GPT-4 and GPT-4o, we used nucleus sampling (temperature = 0.6 and top-p = 0.9) with the standard OpenAI chat template. The LLaMA-3.1-8B and Phi-3-small models were run with greedy decoding and using their respective chat templates.
@@
Appendix C Additional experimental results
C.1 Variance over seeded run results
@@
Note that the same seed is used both for the response generation part and the GPT-4-based evaluation part, as both can be sources of variance in the reported results. Based on our configuration search (see Appendix B.2), some of the response generation models were set to use greedy decoding: LongAlpaca-7B, LongAlpaca-13B, LongChat-7B-v1.5, LongAlign-7B, LLaMA-3.1-8B and Phi-3-small. For these models, the response generation is deterministic and the only source of variance is that of the GPT-4 evaluator.
The results from Table 9 show that the variance across settings is fairly different. In the single-turn ELITR-Bench-QA setting, the standard deviation for all models remain relatively low, even for the models that use nucleus sampling (GPT-3.5, GPT-4, GPT-4o, Vicuna-7B-v1.5, Vicuna-13B-v1.5, LongAlign-13B). However, in the multi-turn settings, we observe an increased standard deviation for those same models overall, in particular for Vicuna-7B-v1.5 and LongAlign-13B. We hypothesize that the sequence of questions asked in the same conversation in the multi-turn setting causes different seeded runs to cumulate errors and slightly diverge along the course of the conversation.
Table 9: Results for the seeded runs on the test set for different ELITR-Bench settings. The reported numbers correspond to the mean score ± sample standard deviation computed over 3 seeds. Boldface numbers correspond to the best performance among proprietary or open models. The results for GPT-3.5 are omitted in the multi-turn setting as the context length exceeded the 16k limit of this model.
@@
LLaMA-3.1-8B 7.83 ± 0.03 7.81 ± 0.02 7.78 ± 0.03
Phi-3-small 7.34 ± 0.05 7.52 ± 0.08 7.38 ± 0.07
C.2 Results on QA/Conv differentiating questions
Similarly with what we observed in Table 2, we notice again a clear difference between the top-performing models – GPT-4/4o, LLaMA-3.1-8B, and Phi-3-small – and the LLaMA-2-based models: the performance of the former is maintained from single-turn to multi-turn, whereas the performance of the latter notably degrades. In contrast with our previous findings that showed little to no difference between the results on ELITR-Bench-QA and ELITR-Bench-Conv for the multi-turn mode, we observe this time that the average score decreases from QA to Conv for LLaMA-2-based models. While the difference is small, this trend is expected as the Conv questions in this subset are more challenging to answer due to the pronominal references. We hypothesize that the opposite trends identified for top-performing models and LLaMA-2-based models might be explained by a ‘snowballing’ effect that causes an error propagation in lower-performing models. It is also possible that GPT-4/4o and the more recent open LLMs (LLaMA-3.1-8B and Phi-3) have been fine-tuned on a larger volume of conversational (i.e., generally multi-turn) data which could imply a greater adaptation to this setting.
C.3 Investigation of a “lost in the middle” effect